OpenStudy (anonymous):

Please Help!!!! Fan and Medal!!!!!! When rolling a number cube, what is the probability of rolling a number greater than 2? A. 1/2 B. 2/3 C. 1/4 D. 5/6

OpenStudy (sidsiddhartha):

when ur rolling a cube there are 6 possible out comes and only 4 numbers that is 3,4,5,6, are greater than 2 so prob=4/6=2/3

OpenStudy (sidsiddhartha):

dont u get it @sammy34daisy